OpenStudy (anonymous):

All answers must be integers! 2+2x+y+z=1 3w-x-2y+z=0 w+2x+y-z=7 w-x+2y-2z=4

OpenStudy (goformit100):

solve them by matrix methad

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but it takes too long and i don't know how to it correctly.

OpenStudy (amriju):

then go for determinants

OpenStudy (anonymous):

?

OpenStudy (amriju):

cramers rule...

OpenStudy (anonymous):

never heard of it our teacher ha not explained that

OpenStudy (unklerhaukus):

\[\left[\begin{array}{cccc} 0&2&1&1\\3&-1&-2&1\\1&2&1&-1\\1&-1&2&-2\end{array}\right|\left.\begin{array} {c}w\\ x\\y\\z\end{array}\right]=\left[\begin{array} {c}-1\\0\\7\\4\end{array}\right]\]